OpenStudy (tgstudios):

Which of the following options results in a graph that shows exponential growth? f(x) = 0.4(3)x f(x) = 3(0.5)x f(x) = 0.8(0.9)x f(x) = 0.9(5)−x

OpenStudy (tgstudios):

I legit have no clue please explain well

OpenStudy (mathmale):

Important: You're discussing exponentiation here. Please use the symbol " ^ " to indicate exponentiation. Example: f(x)=0.4(3)^x or\[ f(x) = 0.4(3)^x\]

OpenStudy (tgstudios):

ok yes they are all exponential

OpenStudy (tgstudios):

k got that sry, Ill remember that

OpenStudy (tgstudios):

sorry to rush but I'm on a time crunch, its 11:43 where I am

OpenStudy (tgstudios):

I am your 600th fan btw @mathmale :D

OpenStudy (mathmale):

Thank you very much! Everything clear to you now, or have you more questions?

OpenStudy (tgstudios):

I still need a little bit of explaining, my guess at this problem is 3(0.5)exponent x

OpenStudy (tgstudios):

everything else seemed smaller

OpenStudy (tgstudios):

I rly need explanation on how to solve it, as I said, my answer is option B

OpenStudy (tgstudios):

@mathmale ban hottiealert101 and close his question up top

OpenStudy (tgstudios):

@jabez177

OpenStudy (tgstudios):

@agent0smith

OpenStudy (tgstudios):

omg thank you, please continue... where did you go?

OpenStudy (tgstudios):

he did it again...

OpenStudy (tgstudios):

plz ban him

OpenStudy (tgstudios):

IM JUST TRYING TO GET SCHOOLWORK DONE!!! THIS IS TERRIBLE!!! @jabez177 @agent0smith @ANYONE!!!

jabez177 (jabez177):

Oh, sick! There's a user with the username "anyone." haha And Mathmale is here. I do not know how to do this... :)

OpenStudy (tgstudios):

yea I just realized... sorry @anyone XD

OpenStudy (tgstudios):

ok I'm done.. :(

OpenStudy (agent0smith):

Delete all your pointless posts, you'll be more likely to get help. If the base (which has the exponent attached) is greater than 1, it's exponential growth.